OpenStudy (anonymous):

Use the elimination method to solve the system of equations. a+6b=-1 and a+3b=-4

OpenStudy (lgbasallote):

subtract the two equations a + 6b = -1 - a + 3b = -4 ======== 0 + 3b = 3 divide both sides by 3 3b/3 = 3/3 b = 1 now substitute this into the first equation a + 6b = -1 a + 6(1) = -1 a + 6 = -1 subtract 6 from both sides a = -1 - 6 combine like terms a = -7 does that help?

OpenStudy (anonymous):

yes thank you very much.

OpenStudy (lgbasallote):

welcome